# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ Abstract
This paper analyzes the statistical properties of constrained portfolio formation in a high dimensional portfolio with a large number of assets. Namely, we consider portfolios with tracking error constraints, portfolios with tracking error jointly with weight (equality or inequality) restrictions, and portfolios with only weight restrictions. Tracking error is the portfolio’s performance measured against a benchmark (an index usually), and weight constraints refers to specific allocation of assets within the portfolio, which often come in the form of regulatory requirement or fund prospectus. We show how these portfolios can be estimated consistently in large dimensions, even when the number of assets is larger than the time span of the portfolio. We also provide rate of convergence results for weights of the constrained portfolio, risk of the constrained portfolio and the Sharpe Ratio of the constrained portfolio. To achieve those results we use a new machine learning technique that merges factor models with nodewise regression in statistics. Simulation results and empirics show very good performance of our method.
Preliminary Draft
@@ 1 Introduction
The finance literature related to large portfolio formation and risk and Sharpe Ratio of the large portfolios has not considered compliance-type constraints in the portfolio in statistical analysis. One of the key issues in empirical finance is large portfolio formation. Mainly, the number of assets in a portfolio may be even larger than the time span of the portfolio. However, a related empirical issue is the constraints on large portfolios in practice. One such constraint is tracking error (TE). Tracking error is defined as the standard deviation between fund and the corresponding benchmark’s returns. The mutual fund reports are required to disclose its performance against a benchmark, and usually this is an index, such as SP500 in US, or sector/theme-specific index such as healthcare sector index. Not only is tracking error important for index funds or Exchange Traded Funds (ETF), it serves as a important indicator for risk management and performance evaluation. The tracking error restriction is studied in the past in the finance literature by Rolll (1992), Jorion (2011).
Another key issue in optimal portfolio formation is weight constraints imposed on the portfolio. This type of restriction can be found in fund’s prospectus, and most of the time fund companies would comply with this weight constraint. The weight restriction means that fund manager can invest only on a certain asset class at most a given percentage. For example, the fund can tell at most 20% of non-US stocks can be in the portfolio. These type of constraints are analyzed in U.S. context by Alexander et al. (2004). Also the regulators can impose restrictions in portfolio formation. Bajeux-Besnainou et al. (2011) tells that in France, bond and money market funds cannot hold more than 10% of stocks. In Germany, stocks cannot exceed 35% of the life insurance portfolio, and in Italy it cannot exceed 20% of the portfolio. The weight constraint is sometimes considered jointly with the tracking errors.
We are interested in the following question: can we construct a portfolio that is theoretically guaranteed to be the optimized one given the aforementioned constraints? One of the key issues in the restricted portfolio optimization is the existence of optimal portfolios, which were found and analyzed in Bajeux-Besnainou et al. (2011). However, to this date, there is no high dimensional statistical analysis of these constrained portfolios. By using the factor model literature as developed in Fan et al. (2011), and merged with residual nodewise regression technique as shown in Caner et al. (2023), we analyze constrained portfolios via residual nodewise regression. Specifically, we analyze out-of-sample risk and Sharpe Ratio, and propose estimators for these metrics and also an estimator for the optimal constrained weights in a high dimensional portfolio. This type of research provides us valuable insight into portfolio analysis. We also obtain rates of convergence of our estimators. It is clear that growing number of factors adversely affect the estimation errors as well as the noise. However, we prove in our theorems that our estimators are consistent even when number of assets in the portfolio grows and is larger than the time span of the portfolio. We allow both number of factors, assets and time span to grow although factors grow slower than time span and number of assets.
There are several papers in the past that analyze portfolios, and some of the metrics such as risk and Sharpe Ratio. For example, Fan et al. (2011), Fan et al. (2013) propose principal orthogonal complement thresholding estimator (POET), which analyzes unconstrained large portfolios with a principal components based estimator. Ledoit and Wolf (2017) propose nonlinear shrinkage and single factor shrinkage estimator to estimate covariance matrix of returns and use in financial metrics in an unconstrained portfolio. There is also maximum Sharpe Ratio estimated sparse regression (MAXSER) by Ao et al. (2019), where they can estimate an unconstrained large portfolio when the number of assets grow slower than the time span of the portfolio, and obtain a mean-variance efficient allocation. Those three papers relied on covariance matrix estimation of assets. Callot et al. (2021) provide nodewise regression estimate of the precision matrix of asset returns and obtains consistency results of the estimators of several financial metrics as variance of the portfolio. Recently Caner et al. (2023) merged factor model literature and machine learning to develop residual nodewise regression estimate of the precision matrix of asset returns, and show its consistency in large portfolios. Caner et al. (2023) only analyze the simple restriction of all assets in a portfolio should add up to 100%. Another strand of literature studies weight-constrained portfolios such as Jagannathan and Ma (2003) and DeMiguel et al. (2009). Du et al. (2023) considered high-dimensional portfolio with cardinality constraints. However, none of the aforementioned studies include the compliance-type weight constraint for certain class of assets mentioned above. Compared to all other statistical-econometrics literature we analyze tracking error and inequality constraints on large portfolios. Our simulation and empirical results show strong performance of our proposed method.
@@
Section 2 shows the model. Section 3 considers tracking error constraint. Section 4 has joint analysis of tracking error and equality weights. Section 5 proposes a new estimator for weights of a large portfolio, as well as other metrics, for joint tracking error and inequality constraints. Section 6 has simulations. Section 7 has empirics. Appendix contains all proofs and also case of only weight constraints scenario.
@@ 2 Model
We start with the following model, for the $j$ th asset ($j=1,\cdots,p$) excess returns at time $t$ ($t=1,\cdots,T$)
@@
where ${\boldsymbol{b}_{j}}$ is a $K\times 1$ vector of factor loadings, and ${\boldsymbol{f}_{t}}\mathrel{\mathop{:}}K\times 1$ vector of common factors to all asset returns, and $u_{j,t}$ is the idiosyncratic error term for the $j$ th asset at time $t$. All the factors are observed, and this model is used by Fan et al. (2011). All the vectors and matrices in the paper are in bold. Define the covariance matrix of errors ${\boldsymbol{u}_{t}}=(u_{1,t},\cdots,u_{j,t},\cdots,u_{p,t})^{\prime}\mathrel{\mathop{:}}p\times 1$ vector, as ${\boldsymbol{\Sigma}_{u}}\mathrel{\mathop{:}}=E{\boldsymbol{u}_{t}}{\boldsymbol{u}_{t}}^{\prime}$. We assume $\{({\boldsymbol{f}_{t}},{\boldsymbol{u}_{t}})\}_{t=1}^{T}$ to be a strictly stationary, ergodic, and strong mixing sequence of random vectors. Also, let ${\cal F}_{-\infty}^{0},{\cal F}_{T}^{\infty}$ be the $\boldsymbol{\Sigma}$- algebras generated by $\{({\boldsymbol{f}_{t}},{\boldsymbol{u}_{t}})\}$, for $-\infty<t\leq 0$, and $T\leq t<\infty$, respectively. Denote the strong mixing coefficient as $\alpha(T)\mathrel{\mathop{:}}=\textnormal{sup}_{{\cal A}\in{\cal F}_{-\infty}^{0},{\cal B}\in{\cal F}_{T}^{\infty}}|\mathbb{P}({\cal A})P({\cal B})-\mathbb{P}({\cal A}\cap{\cal B})|.$ Express the asset return matrix as
@@
with ${\boldsymbol{y}}\mathrel{\mathop{:}}p\times T$ matrix, and ${\boldsymbol{B}}\mathrel{\mathop{:}}p\times K$ factor loadings matrix, and ${\boldsymbol{X}}=(f_{1},\cdots,f_{T})$ as $K\times T$ matrix of factors, with ${\boldsymbol{U}}$ as $p\times T$ error matrix.
We will assume the sparsity of the precision matrix of errors: ${\boldsymbol{\Omega}}\mathrel{\mathop{:}}={\boldsymbol{\Sigma}_{u}^{-1}}$, and let ${\boldsymbol{\Omega}_{j}^{\prime}}$ be the row vector of ${\boldsymbol{\Omega}}$. In ${\boldsymbol{\Omega}_{j}}^{\prime}$ the indices of non-zero elements of that $j$ th row is denoted as $S_{j}$, for $l=1,\cdots,p$
$S_{j}\mathrel{\mathop{:}}=\{l\mathrel{\mathop{:}}\Omega_{j,l}\neq 0\},$
@@
where $\Omega_{j,l}$ be the $j$ th row and $l$ th column element of ${\boldsymbol{\Omega}}$. Let $s_{j}\mathrel{\mathop{:}}=|S_{j}|$ be the cardinality of nonzero elements in row ${\boldsymbol{\Omega}_{j}^{\prime}}$. $s_{j}$ will be nondecreasing in $n$. Denote the maximum number of nonzero elements across each row of ${\boldsymbol{\Omega}}$ as $\bar{s}$, $\bar{s}\mathrel{\mathop{:}}=\max_{1\leq j\leq p}s_{j}$. We use feasible residual based nodewise regression to estimate ${\boldsymbol{\Omega}}$ as shown in Caner et al. (2023). To describe briefly, let ${\bf\hat{b}}_{j}$ denote the $K\times 1$ factor loading estimates by OLS
The residual is
${\bf\hat{u}_{j}}={\bf y_{j}}-{\boldsymbol{X}^{\prime}}{\bf\hat{b}_{j}},$
with ${\boldsymbol{y}_{j}}=(y_{j,1},\cdots,y_{j,t},\cdots,y_{j,T})^{\prime}\mathrel{\mathop{:}}T\times 1$ vector of excess asset return of $j$ th asset across time. Denote ${\boldsymbol{y}_{-j}}$ as $\bf y$ matrix without the $j$ th row ($p-1\times T$), and it can be expressed as
with ${\boldsymbol{B}_{-j}}$ as factor loadings matrix without $j$ th row ($p-1\times K$), and ${\boldsymbol{U}_{-j}}$ as the error matrix without $j$ th row, $j=1,\cdots,p$. Let ${\bf\hat{B}_{-j}}\mathrel{\mathop{:}}(p-1)\times K$ matrix of OLS estimates of factor loadings of ${\boldsymbol{B}_{-j}}$. We have the following residual based matrix which will be used in residual based nodewise regression
with ${\bf\hat{U}_{-j}^{\prime}}\mathrel{\mathop{:}}T\times(p-1)$ matrix, and
which is $K\times(p-1)$ OLS estimates of factor loadings, except for asset $j$, and $T>K$. The residual based feasible nodewise regression depends on the following, for $j=1,\cdots,p$
$\hat{\boldsymbol{\gamma}}_{j}=argmin_{\boldsymbol{\gamma}_{j}\in R^{p-1}}\left[\|{\bf\hat{u}_{j}}-{\bf\hat{U}_{-j}^{\prime}}\boldsymbol{\gamma}_{j}\|_{T}^{2}+2\lambda\|{\boldsymbol{\gamma}_{j}}\|_{1}\right],$ (1)
with $\lambda>0$ as a positive sequence with $\|{\boldsymbol{v}}\|_{T}^{2}$ as the norm with a generic $T\times 1$ vector defined as $\|{\boldsymbol{v}}\|_{T}^{2}\mathrel{\mathop{:}}=\frac{1}{T}\sum_{j=1}^{T}v_{j}^{2}$ To build the estimate for the precision matrix of errors, we define each row of the precision matrix of errors ${\bf\hat{\Omega}}_{j}^{\prime}\mathrel{\mathop{:}}1\times p$, $j=1,\cdots,p$
${\bf\hat{\Omega}_{j}^{\prime}}\mathrel{\mathop{:}}={\bf\hat{C}}_{j}^{\prime}/\hat{\tau}_{j}^{2},$
where ${\bf\hat{C}_{j}^{\prime}}\mathrel{\mathop{:}}1\times p$ row vector with $1$ in the $j$ th cell of that row and the remainder of that row is $-\hat{\boldsymbol{\gamma}}_{j}^{\prime}$. To give an example $\hat{\boldsymbol{C}}_{1}^{\prime}\mathrel{\mathop{:}}=(1,-\hat{\boldsymbol{\gamma}}_{1}^{\prime})\mathrel{\mathop{:}}1\times p$ vector. Also
$\hat{\tau}_{j}^{2}\mathrel{\mathop{:}}={\bf\hat{u}_{j}^{\prime}}({\bf\hat{u}_{j}}-{\bf\hat{U}_{-j}^{\prime}}\hat{\boldsymbol{\gamma}}_{j})/T,$
as in Caner et al. (2023). Stacking all rows of ${\bf\hat{\Omega}_{j}^{\prime}}$, we form ${\bf\hat{\Omega}}$, which is residual based feasible nodewise estimate of the precision matrix of errors. Next, we define the covariance matrix of returns
${\boldsymbol{\Sigma}_{y}}\mathrel{\mathop{:}}={\boldsymbol{B}}({\bf cov\boldsymbol{f}_{t}}){\boldsymbol{B}^{\prime}}+{\boldsymbol{\Sigma}_{u}},$
@@
with ${\boldsymbol{c}\boldsymbol{o}\boldsymbol{v}\boldsymbol{f}_{t}}$ as the covariance matrix of factors ($K\times K$). By Sherman-Morrison-Woodbury formula, with ${\bf\Theta}\mathrel{\mathop{:}}={\bf\Sigma_{y}^{-1}}$
$\boldsymbol{\Theta}\mathrel{\mathop{:}}=\boldsymbol{\Omega}-\boldsymbol{\Omega}\boldsymbol{B}[(\boldsymbol{c}\boldsymbol{o}\boldsymbol{v}\boldsymbol{f}_{t})^{-1}+\boldsymbol{B}^{\prime}\boldsymbol{\Omega}\boldsymbol{B}]^{-1}\boldsymbol{B}^{\prime}\boldsymbol{\Omega},$
and its estimate based on residual based feasible nodewise estimate of errors is:
$\hat{\boldsymbol{\Theta}}\mathrel{\mathop{:}}=\hat{\boldsymbol{\Omega}}-\hat{\boldsymbol{\Omega}}\hat{\boldsymbol{B}}[(\widehat{\boldsymbol{c}\boldsymbol{o}\boldsymbol{v}\boldsymbol{f}_{t}})^{-1}+\hat{\boldsymbol{B}}^{\prime}\hat{\boldsymbol{\Omega}}_{sym}\hat{\boldsymbol{B}}]^{-1}\hat{\boldsymbol{B}}^{\prime}\hat{\boldsymbol{\Omega}},$
with $\widehat{\boldsymbol{c}\boldsymbol{o}\boldsymbol{v}\boldsymbol{f}_{t}}\mathrel{\mathop{:}}T^{-1}\boldsymbol{X}\boldsymbol{X}^{\prime}-T^{-2}\boldsymbol{X}\boldsymbol{1}_{T}\boldsymbol{1}_{T}^{\prime}\boldsymbol{X}^{\prime}$, with $\boldsymbol{1}_{T}$ representing a $T\times 1$ vector of ones, and $\hat{\boldsymbol{\Omega}}_{sym}\mathrel{\mathop{:}}=\frac{1}{2}[\hat{\boldsymbol{\Omega}}+\hat{\boldsymbol{\Omega}}^{\prime}]$, $\hat{\boldsymbol{B}}^{\prime}=(\boldsymbol{X}\boldsymbol{X}^{\prime})^{-1}\boldsymbol{X}\boldsymbol{y}\mathrel{\mathop{:}}K\times p$ matrix.
@@
Before assumptions we make clear an important point, we allow both $p$ (no of assets), $K$ (no of factors) grow when $T$ increases. To save notation we do not subscript $p,K$ with $T$. We also restrict $min(p,T)>K$.

### v2

# Portfolio Analysis in High Dimensions with TE and Weight Constraints

@@ Abstract
This paper explores the statistical properties of forming constrained optimal portfolios within a high-dimensional set of assets. We examine portfolios with tracking error constraints, those with simultaneous tracking error and weight restrictions, and portfolios constrained solely by weight. Tracking error measures portfolio performance against a benchmark (typically an index), while weight constraints determine asset allocation based on regulatory requirements or fund prospectuses. Our approach employs a novel statistical learning technique that integrates factor models with nodewise regression, named the Constrained Residual Nodewise Optimal Weight Regression (CROWN) method. We demonstrate its estimation consistency in large dimensions, even when assets outnumber the portfolio’s time span. Convergence rate results for constrained portfolio weights, risk, and Sharpe Ratio are provided, and simulation and empirical evidence highlight the method’s outstanding performance.
Keywords: Large constrained portfolios; Factor model; Compliance-type constraint; Residual-based nodewise regression.
@@ 1 Introduction
A central concern in empirical portfolio management is the formation of optimal large portfolios, where the number of assets may exceed the time span of the portfolio. An equally significant challenge relates to the practical constraints imposed on large portfolios.
One notable constraint pertains to tracking error (TE from now on), defined as the standard deviation between a portfolio’s returns and those of its corresponding benchmark. Mutual fund reports are mandated to disclose performance against a benchmark, typically an index like the S&P 500 in the U.S., or a sector/theme-specific index such as the healthcare sector index. The significance of TE extends beyond index funds or Exchange Traded Funds (ETFs), as it serves as a crucial indicator for risk management and performance evaluation. The intricacies of TE restrictions have been investigated in previous studies within the finance literature, see, for example, Roll (1992) and Jorion (2003).
Another pivotal aspect of optimal portfolio formation is the imposition of weight constraints. This restriction is typically outlined in a fund’s prospectus, and fund companies routinely adhere to these constraints. Weight restrictions entail that fund managers can allocate only a specified percentage to a particular asset class. For instance, a fund might specify a maximum allocation of 20% to non-U.S. stocks in the portfolio. Sector funds frequently incorporate weight restrictions on specific industries. The analysis of such constraints within the U.S. context is explored by Alexander et al. (2004). Additionally, regulatory bodies have the authority to impose constraints on portfolio formation. In France, for instance, bond and money market funds are generally restricted from holding more than 20% of stocks. Notably, the consideration of weight constraints is sometimes intertwined with TE. These type of constraints are used since they reduce risk by diversification and shrinking the universe of plausible portfolios.
In this paper, our primary focus is on addressing the question: “Can we construct a portfolio that is theoretically guaranteed to be or close to the optimal one given the aforementioned constraints?” The existence of optimal portfolios under constraints has been previously examined (see Bajeux-Besnainou et al. (2011) and Clark et al. (2022)). However, a notable gap persists in the literature regarding a high-dimensional statistical analysis of these constrained portfolios. Drawing on the factor model literature from Fan et al. (2011) and integrating the residual nodewise regression technique as outlined in Caner et al. (2023), we conduct a comprehensive examination of constrained portfolios in a high-dimensional setting. We name our approach the Constrained Residual Nodewise Optimal Weight Regression (CROWN) method and analyze the estimated portfolio’s out-of-sample risk and Sharpe Ratio under TE and weight constraints. We establish rates of convergence for our estimators of the optimal portfolio weights, risks, and Sharpe Ratio, demonstrating their consistency even when the number of assets in the portfolio exceeds the portfolio’s time span in the constrained framework. Our framework accommodates the growth of factors, assets, and time span, with factors growing at a slower rate compared to time span and the number of assets.
There is extensive literature on portfolio analysis that offers in-depth discussions on related measures such as risks and Sharpe Ratios. For instance, Fan et al. (2011), Fan et al. (2013) propose the Principal Orthogonal Complement Thresholding Estimator (POET), which examines unconstrained large portfolios using a principal components-based estimator. Additionally, Ding et al. (2021) further demonstrate the properties of minimum variance portfolio based on POET. Ledoit and Wolf (2017) propose nonlinear shrinkage and single-factor shrinkage estimators to estimate the covariance matrix of returns for use in financial metrics within an unconstrained portfolio. These papers rely on the estimation of the covariance matrix of assets. Cai et al. (2020) explore the estimation of the precision matrix via CLIME (with $l_{1}$ constraint on the elements of the precision matrix) and investigate the potential gains in the performance of the minimum variance portfolio using high-frequency data. There is also the MaXimum Sharpe Ratio Estimated & sparse Regression (MAXSER) proposed by Ao et al. (2019), which enables the consistent estimation of an unconstrained mean-variance efficient large portfolio. Kan et al. (2024) study distributional properties of the in-sample and out-sample Sharpe Ratios of estimated portfolios.
Callot et al. (2021) offer a nodewise regression estimate of the precision matrix of asset returns, demonstrating consistency results for portfolio weights and risk. More recently, Caner et al. (2023) merged the factor model literature with residual-based nodewise regression to develop Sharpe Ratio estimators, showcasing their consistency in large portfolios. Note that our paper is not a simple extension of Caner et al. (2023). Caner et al. (2023) is mainly about the precision matrix of returns estimation and analysis of Sharpe Ratios without weight or TE constraints. Also most of the analysis of Sharpe Ratio is in-sample as in Theorem 4 of Caner et al. (2023). Our paper has three major differences compared to Caner et al. (2023). First, our paper analyzes $l_{1}$ estimation error of portfolio weights and variance of portfolios with constraints in a high dimensional context. Second, we are also interested in out-sample Sharpe Ratio analysis with constraints, which is absent in Caner et al. (2023). The third difference between the two papers is the proof technique difference. Caner et al. (2023) analyzes the maximum out of sample Sharpe Ratio in their section 4.3 (without portfolio weights added to one) and can come up with a consistency result only the number of assets, $p$ is smaller than the time span of the portfolio $T$. Remark 2-Theorem 8 of Caner et al. (2023) explains this issue, and the main reason for this limited result is the proof of their Theorem 8 crucially depends on the maximum eigenvalue of the covariance matrix of asset returns which diverge at $p$ rate. Here in this paper, our results in the unconstrained case (Lemma A.4), as well as constrained cases (Theorems 1-3, Lemma A.3) prove consistency when $p>T$, and our proof technique is different without using a maximal eigenvalue, but instead benefiting from a sup-norm of the covariance matrix of asset returns which diverge at a much slower rate than $p$. Another related paper is Bajeux-Besnainou et al. (2011), where they find a formula for portfolio weights when there are TE and weight constraints. Our paper focuses on high dimensions and estimation of metrics of large portfolios. Compared to their paper, we also develop a data-based way of differentiating between binding and non-binding constraints and suggest how it can be done and estimated in our Theorem 3 here. That is a new result and a new idea.
Another strand of literature studies weight-constrained portfolios as seen in works such as Jagannathan and Ma (2003), DeMiguel et al. (2009), Fan et al. (2012). Du et al. (2023) considered a high-dimensional portfolio with cardinality constraints. In contrast to other statistical-econometrics literature, our paper is the first one to focus on TE and inequality constraints in the context of large portfolios. Our simulation and empirical results demonstrate the strong performance of the proposed method.
@@
Section 2 introduces the model, while Section 3 considers the TE constraints. Section 4 presents a joint analysis of TE and equality weights. In Section 5, a novel estimator for weights in a large portfolio, along with other metrics, is proposed for joint TE and inequality constraints. Section 6 presents a subset of all simulations; the rest is in Appendix. Section 7 offers empirical studies. Appendix includes all proofs and addresses the scenario involving only weight constraints, a comparison with an unconstrained portfolio, and simulations-empirics comparing several methods.
@@ 2 Model
In this section, we start with our factor model and then discuss our estimator for a precision matrix of asset returns. After that, we analyze assumptions. We start with the following model, for the $j$ th asset ($j=1,\cdots,p$) excess returns at time $t$ ($t=1,\cdots,T$)
@@
where $\boldsymbol{b}_{j}$} is a $K\times 1$ vector of factor loadings, and $\boldsymbol{f}_{t}$} is $K\times 1$ vector of common factors to all asset returns, and $u_{j,t}$ is the idiosyncratic error term for the $j$ th asset at time $t$. All the factors are observed, and this model is used by Fan et al. (2011). All the vectors and matrices in the paper are in bold. Define the covariance matrix of errors ${\boldsymbol{u}_{t}}=(u_{1,t},\cdots,u_{j,t},\cdots,u_{p,t})^{\prime}:p\times 1$ vector, as ${\boldsymbol{\Sigma}_{u}}:=E[{\boldsymbol{u}_{t}}{\boldsymbol{u}_{t}}^{\prime}]$. We assume $\{({\boldsymbol{f}_{t}},{\boldsymbol{u}_{t}})\}_{t=1}^{T}$ to be a strictly stationary, ergodic, and strong mixing sequence of random vectors. Also, let ${\cal F}_{-\infty}^{0},{\cal F}_{T}^{\infty}$ be the $\boldsymbol{\Sigma}$- algebras generated by $\{({\boldsymbol{f}_{t}},{\boldsymbol{u}_{t}})\}$, for $\infty<t\leq 0$, and $T\leq t<\infty$, respectively. Denote the strong mixing coefficient as $\alpha(T):=\textnormal{sup}_{{\cal A}\in{\cal F}_{-\infty}^{0},{\cal B}\in{\cal F}_{T}^{\infty}}|P({\cal A})P({\cal B})-P({\cal A}\cap{\cal B})|.$ Express the assets returns matrix as
@@
with ${\boldsymbol{y}}:p\times T$ matrix, and ${\boldsymbol{B}}:p\times K$ factor loadings matrix, and ${\boldsymbol{X}}=(f_{1},\cdots,f_{T})$ as $K\times T$ matrix of factors, with $\boldsymbol{U}$} as $p\times T$ error matrix.
@@
We will assume the sparsity of the precision matrix of errors: ${\boldsymbol{\Omega}}:={\boldsymbol{\Sigma}_{u}^{-1}}$, and let $\boldsymbol{\Omega}_{j}^{\prime}$} be the $j$ th row vector of $\boldsymbol{\Omega}$}. In $\boldsymbol{\Omega}_{j}$$^{\prime}$ the indices of non-zero elements of that $j$ th row is denoted as $S_{j}$, for $l=1,\cdots,p$, $S_{j}:=\{l:\Omega_{j,l}\neq 0\},$ where $\Omega_{j,l}$ be the $j$ th row and $l$ th column element of $\boldsymbol{\Omega}$}. Let $s_{j}:=|S_{j}|$ be the cardinality of nonzero elements in row $\boldsymbol{\Omega}_{j}^{\prime}$}. $s_{j}$ will be nondecreasing in $T$. Denote the maximum number of nonzero elements across each row of $\boldsymbol{\Omega}$} as $\bar{s}$, $\bar{s}:=\max_{1\leq j\leq p}s_{j}$. We use the feasible residual-based nodewise regression to estimate $\boldsymbol{\Omega}$} as shown in Caner et al. (2023). Denote this estimator as $\hat{\boldsymbol{\Omega}}$. A detailed description is in Appendix B.2. $\hat{\boldsymbol{\Omega}}$ is not symmetric. The covariance matrix of returns $\boldsymbol{\Sigma}_{y}$} can be decomposed to
${\boldsymbol{\Sigma}_{y}}={\boldsymbol{B}}(\boldsymbol{\Sigma}_{f}){\boldsymbol{B}^{\prime}}+{\boldsymbol{\Sigma}_{u}},$
@@
with $\boldsymbol{\Sigma}_{f}$ as the covariance matrix of factors ($K\times K$). By Sherman-Morrison-Woodbury formula, with ${\bf\Theta}:={\bf\Sigma_{y}^{-1}}$
$\boldsymbol{\Theta}:=\boldsymbol{\Omega}-\boldsymbol{\Omega}\boldsymbol{B}[\boldsymbol{\Sigma}_{f}^{-1}+\boldsymbol{B}^{\prime}\boldsymbol{\Omega}\boldsymbol{B}]^{-1}\boldsymbol{B}^{\prime}\boldsymbol{\Omega},$
and its estimate based on the feasible residual based nodewise estimate of errors is:
$\hat{\boldsymbol{\Theta}}:=\hat{\boldsymbol{\Omega}}-\hat{\boldsymbol{\Omega}}\hat{\boldsymbol{B}}[\widehat{\boldsymbol{\Sigma}_{f}}^{-1}+\hat{\boldsymbol{B}}^{\prime}\hat{\boldsymbol{\Omega}}_{sym}\hat{\boldsymbol{B}}]^{-1}\hat{\boldsymbol{B}}^{\prime}\hat{\boldsymbol{\Omega}}$, (1)
with $\hat{\boldsymbol{\Omega}}_{sym}:=\frac{1}{2}[\hat{\boldsymbol{\Omega}}+\hat{\boldsymbol{\Omega}}^{\prime}]$, and the estimator of the covariance matrix of factors is defined by $\widehat{\boldsymbol{\Sigma}_{f}}:=T^{-1}\boldsymbol{X}\boldsymbol{X}^{\prime}-T^{-2}\boldsymbol{X}\boldsymbol{1}_{T}\boldsymbol{1}_{T}^{\prime}\boldsymbol{X}^{\prime}$, with $\boldsymbol{1}_{T}$ representing a $T\times 1$ vector of ones. Since we need symmetricity of the square bracketed inverse term in (1) we use $\hat{\boldsymbol{\Omega}}_{sym}$ there, and $\hat{\boldsymbol{B}}^{\prime}=(\boldsymbol{X}\boldsymbol{X}^{\prime})^{-1}\boldsymbol{X}\boldsymbol{y}:K\times p$ matrix. In the other parts of the formula, for (1) symmetricity of $\hat{\boldsymbol{\Theta}}$ is not needed; see Caner et al. (2023) for this point.
@@
Before assumptions, we allow both $p$ (number of assets) and $K$ (number of factors) to grow when $T$ increases. To save notation, we do not subscript $p,K$ with $T$. We also restrict $\min(p,T)>K$.